You are a shareholder in a C corporation. The corporation earns $2.27 per share before taxes. Once it has paid taxes it will distribute the rest of its earnings to you as a are paid? dindend Assume the corporate tax rate is 38% and the personal tax rate on both dividend and non-dividend income is 20% How much is left for you after all taxes The amount that remains is $ per share. (Round to the nearest cent) on
Answer #1

Answer:

Earning per share E=$2.27

Corporate Tax C=38%

Income Tax I= 20%

So amount that remains =E*(1-C)*(1-I)=2.27*(1-38%)*(1-20%)=$1.126 per share